# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v2

@@ Abstract
In recent years, spatio-temporal graph neural networks (GNNs) have attracted considerable interest in the field of time series analysis, due to their ability to capture, at once, dependencies among variables and across time points. The objective of this systematic literature review is hence to provide a comprehensive overview of the various modeling approaches and application domains of GNNs for time series classification and forecasting. A database search was conducted, and 281 journal papers were selected for a detailed examination of the current state-of-the-art in the field. This examination is intended to offer to the reader a comprehensive collection of proposed models, links to related source code, available datasets, benchmark models, and fitting results. All this information is hoped to assist researchers in future studies. To the best of our knowledge, this is the first systematic literature review presenting a detailed comparison of results of current spatio-temporal GNN models applied to different domains. In its final part, this review discusses current limitations and challenges in the application of spatio-temporal GNNs, such as comparability, reproducibility, explainability, poor information capacity, and scalability.
@@ 1 Introduction
In recent years, graph neural networks (GNNs) have emerged as a powerful class of artificial neural network models aimed at processing data that can be represented as graphs [Zheng2024]. GNNs are in fact particularly well-suited for a wide range of practical and engineering applications where data naturally lend themselves to be represented as graph structures, such as applications to transportation networks [review_traffico_jiang], image analysis [GNN_image], classification of molecular structure [GNN_molecule], and natural language processing [GNN_NLP]. The intuitive concept behind the GNN approach is that the nodes of a graph can be put in correspondence to objects or concepts, whereas the edges can be assumed to represent their mutual relationships [Scarselli2009]. Data for these quantities can then be processed at once by a GNN model. GNNs can be used at three different graph levels, related to three distinct classes of problems: at graph level, at edge level, and at node level [review_traffico_con_codici]. In graph-level problems, the goal is to predict a property of a graph based on its entire structure, rather than on individual nodes or edges. For example, a molecule from a sample can be represented as a graph, and global properties of this molecule can be inferred from data coming from its entire structure in relation to the sample. In edge-level problems, the goal is to predict the presence or absence of edges between pairs of nodes. As an example, this task can be used in recommendation systems to predict potential connections between users and items on the basis of past interactions. In node-level problems, the goal is to predict the role or value of each node within a graph, in order to solve either a classification or a regression task. Noticeably, regression tasks can sometimes be put in the form of temporal “autoregressions”, that is, they can include time.
@@
The inclusion of time in GNN modeling has proven to have far-reaching consequences. For this reason, this review examines the available journals’ literature on the specific use of GNNs on time series related tasks, with a particular focus on time series classification and forecasting. The idea behind the use of GNNs for time series problems is that GNNs can be made capable of capturing complex relationships, both inter-variable (connections between different variables within a multivariate series) but also inter-temporal (dependencies between different points in time) at once [review_Alippi]. The inclusion of this capability in GNN models results in a spatio-temporal GNN approach. Within this approach, the spatial dimension is related to the multivariate framework, where nodes typically correspond to different features of a multivariate time series at a given time step $t$, and edges to their relationships, and the temporal dimension is related to the temporal nature of the data. Namely, spatio-temporal GNNs are a class of GNN models designed to handle data in both spatial and temporal dimensions at once. In a spatio-temporal graph, data in each node or edge can evolve over time, and the challenge is then to model their related dynamic interactions over time.
@@
Notwithstanding all the potential interest for such important features, to the best of our knowledge there is a lack of systematic literature reviews (SLRs) on spatio-temporal GNN models for time series applications. Most of the existing surveys are not SLRs, or they are not recent, or only focus on a limited number of application domains and specific issues not necessarily related to time series. For example, recent reviews focus on algorithmic features [cit_rev2, cit_rev3, survey_Lucheroni, survey_Chikwendu], GNN characterization [cit_rev8, review_Alippi], specific fields [cit_rev4], and distributed training [cit_rev5]. Only one SLR on spatio-temporal GNNs was actually recently published, by [survey_spatiotemporal]. However, due to the restrictive search query used in that paper, that SLR covers 52 publications only. In addition, [survey_kumar_spatiotemporal_data] proposed a recent survey on spatio-temporal predictive modeling techniques for different domains. However, the authors approach the topic from a more general perspective, analyzing different machine learning models without placing too much emphasis on spatio-temporal GNNs. As an organized summary of these contributions, a list with a short description of the main reviews on spatio-temporal GNN models is provided in Tab. LABEL:t:tabella_altre_review.
@@
Table 1: Short description of main reviews on spatio-temporal GNN models.
@@
Wu et al. (2020) [cit_rev3] One of the first reviews on the use of GNNs in machine learning and data mining. It introduces a taxonomy that classifies GNN architectures into four categories: recurrent, convolutional, graph autoencoders, and spatial-temporal GNNs. The focus of this review is primarily on early GNN models, rather than applications on time series data.
@@
Zhou et al. (2020) [review_traffico_con_codici] The authors outline a general design pipeline for GNN models, covering key aspects such as the graph structure, pooling techniques, loss function design, and components of the propagation module (including recurrent and convolutional operators, and skip connections). They present some applications of GNNs, including physical systems modeling, chemical analysis, text processing, image classification, recommendation systems, and analysis of other types of data.
@@
Jin et al. (2023) [cit_rev4] Review on the applications of spatio-temporal GNNs to predictive learning in urban computing. In the specific, the authors describe four methods for the construction of pre-defined graphs: topology-based, distance-based, similarity-based and interaction-based approaches. Then, they discuss various design strategies for spatio-temporal GNNs and their combination with some advanced learning frameworks, such as adversarial learning, continuous learning, and physics informed learning.
@@
Li et al. (2023) [li2023graphneuralnetworkspatiotemporal] The authors aim to provide a systematic review of spatio-temporal mining using GNNs. After providing an overview of the graph structure and the spatial and temporal components of GNNs, the authors describe some applications in fields such as transportation, disaster management, environment, power systems, and neuroscience. However, unlike our review, they do not outline the selection process of the papers or include detailed and comprehensive tables that cover datasets, benchmarks and results.
@@
Lin et al. (2023) [cit_rev5] This review is specialized on distributed training of GNNs. It investigates optimization techniques used in this context, including parallel mini-batch generation, dynamic mini-batch allocation, mini-batch transmission pipelining, and parallel aggregation with edge partitioning.
@@
Longa et al. (2023) [cit_rev2] The authors offer an overview of GNNs for temporal graphs. They describe supervised and unsupervised learning tasks on temporal graphs, and present a taxonomy of temporal GNNs that groups the model in snapshot-based and event-based approaches.
@@
Chen et al. (2024) [survey_Lucheroni] One of the first reviews on the use of GNNs for time series modeling. The authors thoroughly describe a wide range of models for time series classification, forecasting, and anomaly detection, based either on recurrent/convolutional GNNs and attentional GNNs. Moreover, they provide a brief comparison of the models’ performances in traffic forecasting and anomaly detection in water treatment.
@@
Chikwendu et al. (2024) [survey_Chikwendu] This survey provides an overview of graph neural architecture search, and explores methods for representing nodes, edges, and sub-graphs, which add context and semantics to graphs. The authors also describe some applications, including computer vision, brain networks, recommendation system, biomedical application, natural language processing, and traffic.
@@
Feng et al. (2024) [feng2024comprehensivesurveydynamicgraph] This survey examines 81 dynamic GNN models. It introduces a novel taxonomy that categorizes GNNs into discrete-time dynamic graphs and continuous-time dynamic graphs based on their structures, features, and dynamic modeling methods. In addition, the authors perform numerical experiments which evaluate 12 significant models on 6 graph datasets, which are not inherently related to time series.
@@
Jin et al. (2024) [review_Alippi] The authors present a comprehensive review of GNNs for time series analysis, covering forecasting, classification, anomaly detection, and imputation. They categorize spatio-temporal GNNs according to their three main components: the spatial module, the temporal module, and the overall model architecture. Although they focus on the same topic as our SLR, they do not provide a detailed comparison of the results of existing models in the literature.
@@
Khemani et al. (2024) [Khemani2024] The review discusses general concepts of GNNs, including architectures, techniques, datasets, applications, and challenges, offering a broad analysis that emphasizes these aspects without focusing specifically on time series analysis. It provides a brief description of papers in the literature, detailing application areas, datasets used, models applied, and performance evaluation.
@@
Kumar et al. (2024) [survey_kumar_spatiotemporal_data] This SLR includes 200 papers from high-impact journals and prominent conferences on spatio-temporal predictive modeling techniques across different domains. Although its structure is similar to that of our review, its focus is broader and not specifically on spatio-temporal GNNs, as it also covers and discusses statistical and non-GNN machine learning models.
Wu et al. (2024) [cit_rev8] This review offers a comprehensive survey of GNN models from the perspective of computation theory. In particular, the authors compare existing GNN models through various perspectives: time, memory, communication, parallelism, and execution.
@@
Zeghina et al. (2024) [survey_spatiotemporal] This SLR includes 52 papers on spatio-temporal GNNs across various application domains. Compared to our SLR, it covers a limited set of papers due to a restricted search query and it does not provide detailed and comprehensive tables of datasets, benchmarks, and model results.
@@
The intention of the SLR presented here is thus to provide a broad and comprehensive overview of the applications of spatio-temporal GNNs for time series classification and forecasting in different areas. Necessarily, this SLR also aims to assess if the current popularity of spatio-temporal GNN models is indeed due to their effectiveness, and to evaluate their efficacy and accuracy in different fields. Additionally, it collects results and benchmarks used in the discussed papers in order to assist researchers in their work by synthesizing the available literature on GNN applications. To the best of our knowledge, this is the first review presenting comprehensive tables with all the results from the various models and benchmarks proposed in a highly fragmented literature. This synthesis and comparison of the results is performed by gathering data from the existing literature, without conducting any numerical experiment on data. Moreover, given the large number of selected papers, it was not feasible to include a description of all the proposed models. To compensate this limitation, we provide a description of the most commonly used benchmarks and the most accurate models within each group. In case, for more detailed descriptions of experimental setups and data pre-processing techniques, the reader is referred to the original papers. It is hoped that this collected and distilled knowledge can anyway become a valuable resource and reference for researchers.
@@
During the development of this review, we formulated two sets of questions, one general and one specific, which are outlined below. The answers to the first set want to provide a general overview of the publications in the field, whereas the answers to the second set focus on specific aspects of the proposed GNN models.
@@ 1. GQ1)
    Trend: What is the temporal trend of the number of publications on GNN models? Is there a growing interest in GNNs?
@@ 7. SQ7)
As the discussion will progress, the answers to both the general and the specific questions will be integrated in it. This will hopefully provide a comprehensive overview of the state-of-the-art of spatio-temporal GNN models. At the same time, a general overview of such GNN models will be assembled by highlighting similarities and differences between the approaches across different fields. In case of need of an even more detailed analysis of a specific discipline, or other aspects of GNN models, the reader is invited to consult the specialized reviews mentioned before [cit_rev1, cit_rev2, cit_rev3, survey_Lucheroni, cit_rev8, cit_rev4, cit_rev5, cit_rev6].
The paper is organized as follows. After this Introduction, Sec. 2 illustrates the methodology used to collect the papers. Sec. 3 provides a general overview of the selected papers. Sec. 4 introduces some fundamental definitions and notions at the basis of spatio-temporal GNN models, and discusses their taxonomy and the determination of the graph structure. Sec. 5 is the core of this review. It explores the diverse domains of application of the spatio-temporal GNN models proposed in the selected papers for the identified domains. A discussion of the findings and the answers to the research questions mentioned above are provided in Sec. 6. Finally, limits, challenges and future research directions are presented in Sec. 7. Sec. 8 concludes the review. A provides a list of all journal papers included in this SLR, together with the year of publication, group they belong to, case study, and nature of the task (e.g., classification or forecasting). All links mentioned in the papers are included in this review. For the reader’s convenience, links that no longer work at the time of submission are marked with a .

### v4

@@ Abstract
In recent years, spatio-temporal graph neural networks (GNNs) have attracted considerable interest in the field of time series analysis, due to their ability to capture, at once, dependencies among variables and across time points. The objective of this systematic literature review is hence to provide a comprehensive overview of the various modeling approaches and application domains of GNNs for time series classification and forecasting. A database search was conducted, and 366 papers were selected for a detailed examination of the current state-of-the-art in the field. This examination is intended to offer to the reader a comprehensive review of proposed models, links to related source code, available datasets, benchmark models, and fitting results. All this information is hoped to assist researchers in their studies. To the best of our knowledge, this is the first and broadest systematic literature review presenting a detailed comparison of results from current spatio-temporal GNN models applied to different domains. In its final part, this review discusses current limitations and challenges in the application of spatio-temporal GNNs, such as comparability, reproducibility, explainability, poor information capacity, and scalability. This paper is complemented by a GitHub repository at https://github.com/FlaGer99/SLR-Spatio-Temporal-GNN.git providing additional interactive tools to further explore the presented findings.
footnotetext: Accepted for publication in Neural Networks
@@ 1 Introduction
In recent years, graph neural networks (GNNs) have emerged as a powerful class of artificial neural network models aimed at processing data that can be represented as graphs [516]. GNNs are in fact particularly well-suited for a wide range of practical and engineering applications where data naturally lend themselves to be represented as graph structures, such as applications to transportation networks [168], image analysis [313], classification of molecular structure [280], and natural language processing [284]. The intuitive concept behind the GNN approach is that the nodes of a graph can be put in correspondence to objects or concepts, whereas the edges can be assumed to represent their mutual relationships [332]. Data for these quantities can then be processed at once by a GNN model. GNNs can be used at three different levels, related to three distinct classes of problems: at graph-level, at edge-level, and at node-level [521]. In graph-level problems, the goal is to predict a property of a graph based on its entire structure, rather than on individual nodes or edges. For example, a molecule from a sample can be represented as a graph, and global properties of this molecule can be inferred from data coming from its entire structure in relation to the sample. In edge-level problems, the goal is to predict the presence or absence of edges between pairs of nodes. As an example, this task can be used in recommendation systems to predict potential connections between users and items on the basis of past interactions. In node-level problems, the goal is to predict the role or value of each node within a graph, in order to solve either a step-wise classification or a regression task. Noticeably, regression tasks can sometimes be put in the form of temporal “autoregressions”, that is, they can include time. When introduced, GNNs were at first applied to static datasets. The inclusion of time in GNN modeling has proven to have far-reaching consequences. Spatio-temporal GNNs have been applied to a wide range of time series tasks, including forecasting, classification, anomaly detection, and data imputation.
@@
This review examines the available literature on the use of spatio-temporal GNNs for time series step-wise classification and future data forecasting, as these are the tasks most commonly addressed in the literature and best illustrate the current developments in spatio-temporal GNNs. The idea behind the use of GNNs for time series problems is that GNNs can be made capable of capturing complex relationships, both inter-variable (connections between different variables within a multivariate series) but also inter-temporal (dependencies between different points in time) at once [176]. The inclusion of this capability in GNN models results in a characteristic spatio-temporal approach. Within this approach, the spatial dimension is related to the multivariate framework, where nodes typically correspond to different features of a multivariate time series at a given time step $t$, and edges to their relationships, and the temporal dimension is related to the temporal nature of the data. Spatio-temporal GNNs can thus be seen as a class of GNN models especially designed to handle data in both spatial and temporal dimensions at once. In a spatio-temporal graph, data in each node or edge can evolve over time, and the challenge is then to model their related dynamic interactions over time.
In the broader landscape of time series modeling, a variety of other approaches have been developed alongside spatio-temporal GNNs. Early statistical frameworks, such as ARIMA, and, more prominently, neural networks designed for sequential data, including RNNs, LSTMs, and GRUs, have been for a long time the standard solution for modeling temporal data. These architectures effectively capture short and medium term time dependencies, but struggle with very long sequences and with explicitly modeling interactions among multiple variables. The attention mechanism and the Transformer architecture, originally developed for handling long, static word sequences in Large Language Models (LLMs), were later adopted to better address long term dependencies in time series [494]. Building on these architectural advances, and supported by the availability of modern GPU hardware, many experiments have been conducted. However, it remains unclear whether these highly parallelizable architectures are truly suitable for time series tasks [367]. The development of LLMs has also stimulated a further line of research in time series modeling. LLMs, trained on massive text datasets for capturing general patterns in language, can then be adapted to a variety of downstream task, that is, as foundational models. Inspired by this approach, researchers have begun developing time series foundation models (TSFMs), designed to learn general representations from diverse temporal datasets [247]. These models, which do not necessarily use LLMs, are trained on large collections of time series to enable them to generalize effectively to new sequences.
All these approaches share the ability to model the dynamics of temporal data, and are inherently sequential models. What differentiates spatio-temporal GNNs from conventional machine learning, LLMs applied to dynamics, and TSFMs is how they handle the mutual relationships among features in multivariate time series. In TSFMs, these relationships are supposed to be learned end-to-end. In contrast, spatio-temporal GNNs explicitly prioritize these relationships. In the case of GNN models based on pre-defined static adjacency matrices, the inter-variable relationships are effectively “forced” into the model, like handcrafted features (which TSFMs could not be able to learn). As for GNNs using self-learned adjacency matrices, these more recent models are naturally biased toward explicitly identifying these relationships, offering a clearly interpretable representation of inter-dependencies.
@@
Notwithstanding all the potential interest for such important features, to the best of our knowledge there is a lack of systematic literature reviews (SLRs) on spatio-temporal GNN models for time series applications. Most of the existing surveys are not SLRs, or they are not recent, or only focus on a limited number of application domains and specific issues not necessarily related to time series. A much broader view on the subject is nowadays necessary. For example, recent reviews focus on algorithmic features [273, 423, 35, 55], GNN characterization [415, 176], specific fields [174, 78], and distributed training [253]. Only two SLR on spatio-temporal GNNs were actually recently published, in [467] and [78]. However, Ref. [467] uses a restrictive search query that limits the SLR to only 52 publications, whereas Ref. [78] examines a broader collection of papers, but it is only focused on traffic flow prediction. In addition, [197] is a recent survey on spatio-temporal predictive modeling techniques for different domains. However, the authors approach the topic from a more general perspective, analyzing different machine learning models without placing too much emphasis on spatio-temporal GNNs. As an organized summary of these contributions, a list with a short description of the main reviews on spatio-temporal GNN models is provided in Tab. LABEL:t:tabella_altre_review.
@@
Table 1: Short description of main reviews on spatio-temporal GNN models. At the time of writing this SLR, articles marked with a $\diamond$ were only available as preprints.
@@
Wu et al. (2020) [423] One of the first reviews on the use of GNNs in machine learning and data mining. It introduces a taxonomy that classifies GNN architectures into four categories: recurrent, convolutional, graph autoencoders, and spatial-temporal GNNs. The focus of this review is primarily on early GNN models, rather than applications on time series data.
@@
Zhou et al. (2020) [521] The authors outline a general design pipeline for GNN models, covering key aspects such as the graph structure, pooling techniques, loss function design, and components of the propagation module (including recurrent and convolutional operators, and skip connections). They present some applications of GNNs, including physical systems modeling, chemical analysis, text processing, image classification, recommendation systems, and analysis of other types of data.
@@
Jin et al. (2023) [174] Review on the applications of spatio-temporal GNNs to predictive learning in urban computing. In the specific, the authors describe four methods for the construction of pre-defined graphs: topology-based, distance-based, similarity-based and interaction-based approaches. Then, they discuss various design strategies for spatio-temporal GNNs and their combination with some advanced learning frameworks, such as adversarial learning, continuous learning, and physics informed learning.
@@
Li et al. (2023) [236] The authors aim to provide a systematic review of spatio-temporal mining using GNNs. After providing an overview of the graph structure and the spatial and temporal components of GNNs, the authors describe some applications in fields such as transportation, disaster management, environment, power systems, and neuroscience. However, unlike our review, they do not outline the selection process of the papers or include detailed and comprehensive tables that cover datasets, benchmarks and results.
@@
Lin et al. (2023) [253] This review is specialized on distributed training of GNNs. It investigates optimization techniques used in this context, including parallel mini-batch generation, dynamic mini-batch allocation, mini-batch transmission pipelining, and parallel aggregation with edge partitioning.
@@
Longa et al. (2023) [273] The authors offer an overview of GNNs for temporal graphs. They describe supervised and unsupervised learning tasks on temporal graphs, and present a taxonomy of temporal GNNs that groups the model in snapshot-based and event-based approaches.
@@
Chen et al. (2024) [35] One of the first reviews on the use of GNNs for time series modeling. The authors thoroughly describe a wide range of models for time series classification, forecasting, and anomaly detection, based either on recurrent/convolutional GNNs and attentional GNNs. Moreover, they provide a brief comparison of the models’ performances in traffic forecasting and anomaly detection in water treatment.
@@
Chikwendu et al. (2024) [55] This survey provides an overview of graph neural architecture search, and explores methods for representing nodes, edges, and sub-graphs, which add context and semantics to graphs. The authors also describe some applications, including computer vision, brain networks, recommendation system, biomedical application, natural language processing, and traffic.
@@
Feng et al. (2024) [90] This survey examines 81 dynamic GNN models. It introduces a novel taxonomy that categorizes GNNs into discrete-time dynamic graphs and continuous-time dynamic graphs based on their structures, features, and dynamic modeling methods. In addition, the authors perform numerical experiments which evaluate 12 significant models on 6 graph datasets, which are not inherently related to time series.
@@
Jin et al. (2024) [176] The authors present a comprehensive review of GNNs for time series analysis, covering forecasting, classification, anomaly detection, and imputation. They categorize spatio-temporal GNNs according to their three main components: the spatial module, the temporal module, and the overall model architecture. Although they focus on the same topic as our SLR, they do not provide a detailed comparison of the results of existing models in the literature.
@@
Khemani et al. (2024) [182] The review discusses general concepts of GNNs, including architectures, techniques, datasets, applications, and challenges, offering a broad analysis that emphasizes these aspects without focusing specifically on time series analysis. It provides a brief description of papers in the literature, detailing application areas, datasets used, models applied, and performance evaluation.
@@
Kumar et al. (2024) [197] This SLR includes 200 papers from high-impact journals and prominent conferences on spatio-temporal predictive modeling techniques across different domains. Although its structure is similar to that of our review, its focus is broader and not specifically on spatio-temporal GNNs, as it also covers and discusses statistical and non-GNN machine learning models.
@@
Zeghina et al. (2024) [467] This SLR includes 52 papers on spatio-temporal GNNs across various application domains. Compared to our SLR, it covers a limited set of papers due to a restricted search query and it does not provide detailed and comprehensive tables of datasets, benchmarks, and model results.
Capone et al. (2025) [28] This review provides an overview of recent spatio-temporal GNN models and classifies them as time-then-space (TTS), space-then-time (STT), and combined space-and-time (S&T) models. Compared to our SLR, this review is not systematic, it includes a limited number of papers, and it does not provide detailed quantitative comparisons of model performance across datasets.
Fan et al. (2025) [78] This SLR covers 86 studies on the use of GNNs for traffic flow prediction. It examines data processing, model deployment, graph construction methods, and network architectures, discussing their respective strengths and weaknesses. However, unlike our review, its focus is on a single application domain.
Wu et al. (2025) [415] This review offers a comprehensive survey of GNN models from the perspective of computation theory. In particular, the authors compare existing GNN models through various perspectives: time, memory, communication, parallelism, and execution.
While most existing reviews focus on specific subfields, the present work takes a unique broader perspective. Its aim is to systematically collect and synthesize a wide range of publications on spatio-temporal GNNs for time series tasks, covering multiple application domains and methodological approaches. As a result, the number of articles included in this survey is larger than in more narrowly focused reviews. Although this approach results in a longer text, we hope it is able to provide a unique overview of the field, capturing both the diversity of applications and recent developments that cannot be fully appreciated in more specialized surveys.
@@
The intention of the SLR presented here is therefore to: (i) offer a wide-ranging overview of spatio-temporal GNNs for time series classification and forecasting in different domains, (ii) assess whether their popularity stems from actual effectiveness, (iii) evaluate their efficacy and accuracy in different fields, (iv) collect results and benchmarks from the literature on GNN applications to support researchers in their work. In addition, to the best of our knowledge, this is the first review presenting comprehensive tables with all the results from the various models and benchmarks proposed in a highly fragmented literature. The tables presented in the paper are also available in the accompanying GitHub repository in an interactive format, enabling users to sort, filter, and explore the results of the various models. This synthesis and comparison of the results is performed by gathering data from the existing literature, without conducting any numerical experiment on data. Moreover, given the large number of selected papers, it was not feasible to include a description of all the proposed models. To compensate this limitation, we provide a description of the most commonly used benchmarks and the most accurate models within each group. In case, for more detailed descriptions of experimental setups and data pre-processing techniques, the reader is referred to the original papers. It is hoped that this collected and distilled knowledge can anyway become a valuable resource and reference for researchers.
@@
During the development of this review, we formulated two sets of questions, one general and one specific, which are outlined below. The answers to the first set want to provide a general overview of the publications in the field, whereas the answers to the second set focus on specific aspects of the proposed spatio-temporal GNN models.
@@ 1. GQ1)
    Trend: What is the temporal trend of the number of publications on spatio-temporal GNN models? Is there a growing interest in GNNs?
@@ 4. GQ4)
    Conferences: What are the most prominent conferences where GNN research is presented?
  5. GQ5)
@@ 7. SQ7)
As the discussion will progress, the answers to both the general and the specific questions will be integrated in it. This will hopefully provide a comprehensive overview of the state-of-the-art of spatio-temporal GNN models. At the same time, a general overview of such GNN models will be assembled by highlighting similarities and differences between the approaches across different fields. In case of need of an even more detailed analysis of a specific discipline, or other aspects of GNN models, the reader is invited to consult the specialized reviews mentioned before [14, 273, 423, 35, 415, 174, 253, 339, 78, 521, 236, 55, 90, 176, 182, 197, 467, 28].
The paper is organized as follows. After this Introduction, Sec. 2 illustrates the methodology used to collect the papers. Sec. 3 provides a general overview of the selected papers. Sec. 4 introduces some fundamental definitions and notions at the basis of spatio-temporal GNN models, and discusses their taxonomy and the determination of the graph structure. Sec. 5 is the core of this review. It explores the diverse domains of application of the spatio-temporal GNN models proposed in the selected papers for the identified domains. A discussion of the findings and the answers to the research questions mentioned above are provided in Sec. 6. Finally, limits, challenges and future research directions are presented in Sec. 7. Sec. 8 concludes the review. Appendix A provides a list of all papers included in this SLR, together with the year of publication, group they belong to, case study, and nature of the task (e.g., classification or forecasting). All links mentioned in the papers are included in this review. For the reader’s convenience, links that no longer work at the time of submission are marked with a . The overall structure of the paper is visually summarized in Fig. 1.
Figure 1: Overview of the structure of the SLR.
To maximize the accessibility and usability of the collected information, the paper is supported by a GitHub repository, available at https://github.com/FlaGer99/SLR-Spatio-Temporal-GNN.git, which complements the review with interactive resources. Beyond listing the 366 reviewed papers organized by groups, the repository offers interactive tables for sorting and filtering models by any metric, dynamic plots to visualize performance and trends, and tools to explore top-performing models and most common benchmarks. These features make it easy for readers to analyze, compare, and gain deeper insights from the collected data. We strongly encourage exploring the repository as an interactive complement to this paper, especially when examining the extensive and complex tables comparing model results, as it allows readers to easily sort and visualize the data.